970 ro Solve: 92 0.06(92-+) (450+6+ -0.05+2e dt FVE Sa 365
Answer #1

Solve F.V. 92 (450 + 6t-o.ost2) e 0.06 (92-t) 365 dt 92 let Is 92 0.06 (92-t) 0.06(92-t) 365 at = -365 0.06 * ) 365 O I -365F.V. – 450 I +6 I2 -0.05 I3 450 ( 92.699191) +6(4,274, 911)-0.05(202,524.63) = 41,714.64 + 25,649.46 – 13,126.23 = 54231.87 T